1. According to the Michaelis-Menten equation, when is the reaction velocity at half the maximal value (vo = 0.5 Vmax)

A. when [S] = 0.5 Km

B. when [S] = 1.0 Km

C. when vo = kcat [E]

D. when vo = kcat / Km

2. The value of Vmax depends on ...

A. k1 and k-1

B. k1 and k2

C. k1 only

D. k2 only

3. When is Km approximately equal to Kd?

A. when k1 is much less than k-1

B. when k2 is much less than k-1

C. when k-1 is much less than k1

D. when k-1 is much less than k2
Answer #1

Question 1

Ans : Option B

\odot When [S] = 1.0 [Km]

Reason :

The Michaelis Menten Kinetic Equation:

Vo = (Vmax [S]) / (Km + [S]) ;

When [S] << Km,

Vo = Vmax [S] / Km

When [S] >> Km,

Vo = Vmax [S] / [S]

Vo = Vmax

When [S] = Km

Vo = Vmax Km / (Km + Km)

Vo = Vmax Km / 2 Km

Vo = Vmax / 2

Vo = 0.5 Vmax

Thus [S] = 1.0 Km at which reaction is half its maximal value.

Question 2

Ans : Option D

\odot k2 only

Reason : Vmax is the product of catalytic constant on enzyme concentration.

That is,

Vmax = kcat [ET]

kcat = catalytic constant = turnover number

Here k2 = kcat

The higher the k2 is, the more substrates get turned over in one second. Therefore Vmax depends on k2 only.

Question 3

Ans : Option B

\odot When k2 is much less than k-1

Reason : The QSSA (quasi-steady-state approximation) defines the Km value as:

Km = k-1 + k2 / k1

where,

k1 is the rate constant

k-1 is the unproductive dissociation rate constant of ES – complex

k2 = kcat is the catalytic rate constant or turnover number

When k2 << k-1,

Km = k-1 + k2 / k1

Km ≈ k-1 / k1

k-1 / k1 = Kd

Therefore ,

Km ≈ Kd